OpenStudy (abmon98):

i)find the first 3 terms in ascending powers of x in the expansion (1+x)^5 the coefficient of x^2 in the expansion of (1+(px+x^2))^5 is 95 use your answer in i to find the value of the constant p

OpenStudy (abmon98):

(1+x)^5=1+5x+10x^2

OpenStudy (abmon98):

@ganeshie8

OpenStudy (abmon98):

how can i find the value of constant p

ganeshie8 (ganeshie8):

(1+(px+x^2))^5 = 1 + 5(px+x^2) + 10(px+x^2)^2 + .. right ?

OpenStudy (abmon98):

shoudnt it be 1+5(px)^4*x^2)

OpenStudy (abmon98):

oh you are right

OpenStudy (abmon98):

sorry for that

OpenStudy (abmon98):

1+5px+5x^2+10p^2x^2+20px^2+10x^4

OpenStudy (abmon98):

20px^3

OpenStudy (abmon98):

10p^2+5=95 p>0 p=3

OpenStudy (kainui):

What does "first terms" mean? To me, there is no difference between x^5 or 1 as being the first term.

OpenStudy (abmon98):

first 3 terms of a bionomial expansion from 1^5 on towards(5,5) (x)^5